Circular No. 8166

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

V2573 OPHIUCHI = NOVA OPHIUCHI 2003 On July 17, Akira Takao, Kitakyushu, Japan, reported the discovery

of a possible nova on unfiltered CCD images taken with a 120-mm f/4 telephoto lens, providing a position of α = 17^{h}19^{m}14^{s}.4, δ = −27^{o}22′37′′ (equinox 2000.0; uncertainty ± 20′′), and the following approximate magnitudes: 2001 Apr. 18.806 UT, [12.6; 2002 Mar. 18.808, [11.6; May 12.647, [11.6; 2003 Apr. 18.812, [11.6; July 10.598, 11.4; 16.523, 11.2. L. A. G. Monard, Pretoria, S. Africa, reported that the object appeared at mag 10.4 on an unfiltered CCD image taken on July 17.821, providing the following precise position end figures: 14^{s}.10, 35′′.4 (uncertainty ± 0′′.2); Monard adds that nothing is visible at this position on a red Digitized Sky Survey image. V. Tabur, Wanniassa, Australian Capital Territory, reports an independent discovery of this variable star on unfiltered CCD images obtained with a 140-mm f/2.8 telephoto lens, explaining that he initially dismissed the object as a less-urgent variable star due to its long presence on his past images, as shown by the following magnitude estimates: 2003 Mar. 21.600, 12.9; Apr. 5.652, 13.1; 29.579, 13.2; May 4.566, 13.1; 10.538, 13.1; 22.667, 13.0; 29.595, 12.6; June 16.534, 10.1; July 5.574, 9.8. (Tabur provides position end figures 14^{s}.2, 34′′, uncertainty ± 5′′). T. Kato, Kyoto University, reports that ASAS-3 images show the new object in outburst on June 15, with a peak of V = 10.6 reached on June 26.

M. Della Valle, Arcetri Astrophysical Observatory, Florence; and E. Mason, L. Pasquini, and J. Prichard, European Southern Observatory, write: “Preliminary analysis of a spectrum (range 380–900 nm, resolution 48000), obtained on July 18.2 UT at La Silla with the 2.2-m telescope (+ Fiber-fed Extended Range Optical Spectrograph), shows the new object to be a nova caught during its early decline. The spectrum is dominated by Hα, O I (mult. 1, 4; measured at 778.0 and 844.0 nm), and Ca II (mult. 2; measured at 850, 855, and 866 nm). Hα, Hβ, and Na I lines are flanked by multiple P-Cyg profiles whose minimum absorptions give maximum expansion velocities of ∼ 1700, 1900, and 1500 km/s, respectively. Also prominent are emission lines due to Fe II (mult. 42, 48, 49, 74). In particular, we note that the 493-nm line (mult. 42) is as strong as Hβ. [O I] at 630 and 636.4 nm and He I at 667.8 nm are also detected. The equivalent widths of the Na D interstellar lines (0.085 and 0.07 nm) and Ca II H and K (0.04 and 0.06 nm) may suggest that this nova suffers from moderate absorption.”

N. N. Samus, Institute of Astronomy, Moscow, informs us that the designation V2573 Oph has been given to this object.

2003 July 27  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green